40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 2578     Entry time: Wed Oct 28 18:10:27 2020
Author: aaron 
Type: Noise Hunting 
Category: Laser 
Subject:  

Lab entry

in: Wed Oct 28 18:11:00 2020
out: Wed Oct 28 21:05:58 2020
 

E laser noise

  • Moved the Newfocus 1611 from W path cryo cantilevers transmission, to just after the PSOMA beam launch.
    • nominal AC gain: 700 V/A
    • nominal DC gain: 10 kV/A
    • output impedance: 50 Ohm
    • Max input power: 1 mW DC (10 mW CW)
  • Measuring the following spectra across the PD band (~30kHz-1GHz) by connecting the 1611 AC output directly to an HP 8560E (50 Ohm input). I'm monitoring DC with TDS3024B oscilloscope (1 MOhm input).
    • dark current
    • free running E path laser intensity noise

Optical layout: beam launch -> lambda/2 -> steering mirrors -> lens 1 -> ND 0.6 -> lens 2 -> PD 1611

There is only a 12 VDC power supply compatible with the 1611 power port, but the PD requires +- 15V. Surely there's one somewhere. Perhaps this is why I observe only -6 V on the DC mon with 1mW input power at 1550nm (checked against the Thorlabs S122C; I expected -10V). Maybe the beam is too large.

ELOG V3.1.3-